Case 15.4: Sapphire Coffee-Part 2 This case is an extension of the Sapphire Coffee case from Chapter 14. Case the issue would be: Is the $5.00 increase in sales for each one square foot increase in store size adequate and if not through windowwing the location of the store and whether it had a drive w help explain sales. A stand alone analysis would start with the output from the case in Chapter 14 with a discussion that the low R-square value and relatively high confidence interval would justify looking for additional independent variables. This would lead to dummy variables: d1 No 0 Yes 1 Near a college d2 Drive through window No 0 Yes The Excel output with the dummy variables is SUMMARY OUTPUT Regression Statistics 0.852547827 0.726837797 Multiple R Adjusted R Square0.71011358 Standard Error Observations 252.0293585 ANOVA df MS Regressiorn Residual 3 8281618.949 2760540 43.4602 7.58205E-14 49 3112421.08 63518.8 52 11394040.03 Coefficients Standard Error tStat P-value Lower 95% Upper 95% 61.063116 273.490089 5.342289 2.37E-06 911.4645391 2010.662 14 -109.362 314.3532146 105.6491529 2.975445 0.004533 102.0434219 526.663 Intercept Store Size (Sq. F) 3.25633342 0289857069 11.234273.685-15 2673844106 3 838823 8452 84.63282732 -1.2922 0.202348 -279.4387761 60.71309 College Nearby Drive Thru Adding the dummy variables increased the R-square value of the model, but knowing if there is a college near-by is not significant. Having a drive through window seems to significantly increase sales but how much is somewhat questionable given the wide confidence interval.
32. Gillians Restaurant has an ice-cream counter where it sells two main pro ducts, ice cream and n yogurt, each in a variety of flavors. The restaurant makes one order for ice cream and yogurt each week, and the store has enough freezer space for 115 gallons total o oroze yogurt costs $0.75 and a gallon of ice cream costs $0.93.and the restaurant each week for these products. The manager estimates that each week the restaurant east twice as much ice cream as frozen yogurt. Profit per gallon of ice cream is $4.15 and profit per gallon of yogurt is $3.60. e a linear programming model for this problenm b. Solve this model by using graphical analysis.
Answer #1

Blue line is : x-2y = 0

Green Line : 0.75x+0.9y = 90

Red Line : x+y = 115